# Eaman Jahani

### v4

@@ D.1 Task Performance and ATEs
The top pane of Figure 1B compares the average performance across models and attempt numbers (also referred to as iterations). It shows the average cosine similarity score stratified by the reference image. A notable feature in this figure is the performance dip during the second recreation attempt across both treatment arms. This is likely due to participants’ initial misunderstanding of the model’s “memoryless” nature. Participants failed to recognize that context from previous prompts was not carried over to new iterations. We observed numerous prompts in the second iteration across users that explicitly referenced the first prompt, a behavior that rarely occurred in subsequent attempts. However, from the third prompt onward, participants appeared to grasp the independence of each attempt, as evidenced by a marked decrease in cross-prompt references and a corresponding rebound in performance.
@@
Next, the bottom pane of Figure 1B shows the average treatment effect (ATE) per iteration, which is the difference between the stratified averages of DALL-E 3 and DALL-E 2 in the top pane. [^4] In Section D, when we refer to “DALL-E 3”, we mean “DALL-E 3 (Verbatim)” unless otherwise specified. To test the widening impact of using DALL-E 3 on performance relative to DALL-E 2, we also run the following fixed effects linear model with participant-level ($i$) clustered standard errors where iteration is treated as a numeric variable:
@@
$Y_{i,n,t}=\beta_{0}+\beta_{1}\text{iteration}+\beta_{2}\mathbb{I}[\text{dalleVersion = 3}]_{i}+\beta_{3}\text{iteration}\times\mathbb{I}[\text{dalleVersion = 3}]_{i}+\gamma_{t}+\epsilon_{i,n,t}$ (6)
@@ D.2 Prompt Characteristics
Figure 1C compares the prompt length and prompt similarity of the two models. To generate these results, we first remove any prompt that does not constitute a good-faith attempt according to the sample construction procedure detailed in Section B.B.4. The prompt length is the average number of words per model and iteration stratified by the reference image. The prompt similarity is the average cosine similarity between all consecutive pairs of user prompts, which are both determined to be valid attempts, stratified by the reference image (see Section B.B.5.3 for details on similarity calculations). The color scale in Figure 1C shows the stratified average similarity to the previous prompt across all users per each model. We find that superior model users write prompts that, on average, have $\beta=0.0184$ higher in cosine similarity to their previous prompts using cluster robust standard errors at the participant level ($p=0.0236$).
@@
Comparing the aggregate similarity of all attempts made by a given participant, we also find the prompts from DALL-E 3 participants were, on average, more similar than the prompts of the inferior model participants. For this analysis, we use the dispersion around the centroid in the prompt embedding space, explained in Section B.B.5.3, as the dependent variable. When we average across all participants by model, we find that the average distance of prompts written by superior model participants to their centroid is $\beta=0.0191$ smaller than inferior users ($p=0.0083$). Standard errors are cluster-robust at the participant level.
D.3 ATE Decomposition
@@
Figure 2 in the main text decomposes the ATE into the model and prompting effects. This decomposition is conceptually similar to a simple mediation analysis, with an important difference being that we can observe counterfactual outcomes (e.g., prompting the superior model as if it is the inferior model). This is not typically the case in mediation analysis, and makes causal identification rely on fewer assumptions. To obtain counterfactual outcomes, we fed or “replayed” the participant prompts when interacting with one model (e.g., inferior) on another model (e.g., superior). The notation (prompt, model) specifies which treatment arm the prompts were written under and which model was used in the replay. For example, (2,3) indicates replaying prompts written under DALL-E 2 on DALL-E 3. To be clear, (2,2) and (3,3) correspond to the original observed treatment arms, while (2,3) and (3,2) are the counterfactual outcomes of interest. [^5] To avoid problems with model drift, we regenerated images for all four possible combinations at the same time and used these images for all analyses in the main text.
@@
The left-most point in Figure 2 corresponds to the average CLIP cosine similarity to the target image of (2,2). To make the interpretation of the results clearer, we have subtracted this quantity from all average quality scores and added a dashed line throughout. The second point from the left corresponds to average similarity to the target of (2,3), the third point from the left to (3,3), and the rightmost points to (3,2). All average similarity scores are stratified by iteration and reference image, and the standard errors are bootstrapped and cluster-robust at the participant level. The model effect, as shown by the red braces in Figure 2, corresponds to the average increase in quality of (2,3) relative to (2,2). In the terminology of mediation analysis, the model effect would be referred to as the direct effect. The prompting effect, as shown by the blue braces in Figure 2, corresponds to the average increase in quality of (2,3) relative to (3,3). In the terminology of mediation analysis, the prompting effect would be referred to as the indirect effect. We can also test the difference in average quality between (3,3) and (3,2), as well as the difference between (3,2) and (2,2). Both of these differences are visible in Figure 2; the second is small and not statistically significant.
@@
The standard errors in Figure 2 correspond to the uncertainty around the estimated average score for each of the four replay conditions. These uncertainty estimates are insufficient for exact inference on the direct, indirect, and treatment effects. The statistics and significance values report in the main text, which correspond to such effects (i.e., the difference between average estimates in two conditions) are obtained using a two-way (iteration and target image) fixed effect model with the effect type as the main independent variable:
@@
Appendix E Robustness Checks
E.1 DreamSim-based Analysis
As discussed in Section B, we repeat all main-text analyses with DreamSim. The results of these analyses are reported below.
E.1.1 Overall ATEs
@@
E.1.2 Figure 1
We reran the regression in Section D.D.1 with $Y_{i,n,t}$ representing the DreamSim outcome instead of cosine similarity:
$Y_{i,n,t}=\beta_{0}+\beta_{1}\text{iteration}+\beta_{2}\mathbb{I}[\text{dalleVersion = 3}]_{i}+\beta_{3}\text{iteration}\times\mathbb{I}[\text{dalleVersion = 3}]_{i}+\gamma_{t}+\epsilon_{i,n,t}$ (8)
@@ E.1.3 Figure 2
E.2 Z-score-based Analysis
As we discuss in Section C.2, we repeat all main-text analyses with the within-image-attempt Z-score of CLIP-cosine similarity to better account for variation between images and attempts. When comparing across attempts, the Z-score is computed within each image, as mentioned in section C.2. The results hold across the board and sometimes, the differences between the superior and inferior model are even starker than when using cosine similarity.
E.2.1 Overall ATEs
@@ E.2.2 Figure 1
$\text{ZScore}_{i,n}=\beta_{0}+\beta_{1}\text{iteration}+\beta_{2}\mathbb{I}[\text{dalleVersion = 3}]_{i}+\beta_{3}\text{iteration}\times\mathbb{I}[\text{dalleVersion = 3}]_{i}+\epsilon_{i,n}$ (9)
@@ E.2.3 Figure 2
Appendix F Pre-registration
Prior to our experiment, we pre-registered a number of hypotheses and a pre-analysis plan. This pre-registration is deposited at OSF at the following URL: https://osf.io/ejbtp. The main text of this paper contains a subset of our pre-registered analysis, as well as complementary exploratory analyses that are also mentioned in our pre-registration. We chose to present this subset of our pre-registered analyses because we believe this subset constitutes an important and timely set of results best-suited to a short-form paper.
Below, we provide a high-level description of all of our pre-pregistered analyses, along with a description of the results of those analyses. In our pre-registration, we declared the intent to conduct each of our analyses using six666see section F.2 on deviations from the pre-registration for the remaining two pre-registered outcome variables. possible outcome variables, all of which are different transformations of the same underlying data: CLIP embedding cosine similarity and DreamSim, both of which rescaled in three ways:
  • No rescaling: The outcome variable is used as-is.
  • Z-score rescaling: We rescale the outcome variable into a Z-score according to the procedure describe in these supplementary materials.
  • Percentile rank rescaling: We rescale the outcome variable into a percentile rank. This is done by calculating the percentile rank of a given prompt relative to all other prompts submitted for the relevant target image.
We intend to complete each pre-registered analysis with each of these eight possible outcome variables. Although we have not yet done so, we have no reason to believe that our results will not be robust to these different transformations of our data. However, we did not want to delay the publication of our preprint until all of these analyses were complete. Thus, we are posting a preprint with our pre-registered analyses complete for only a subset of these pre-registered outcomes. As we complete more of these analyses, we will update this supplement. We also intend to post full results, data, and replication code online shortly.
F.1 Hypotheses and Results
Below, we list the hypotheses precisely as they are written in our pre-registration document. For each hypothesis, we also describe our results.
@@ F.1 Hypotheses and Results
Analysis approach:
For this hypothesis, we conducted multiple ANOVA tests, one for each of the demographic variables, against the relevant outcome variable and treatment variable (the model) as the covariate. As a robustness check, we repeated the same procedure using the Kruskal-Wallis U test. To adjust for multiple testing, we used the Benjamini-Hochberg adjustment with a false discovery rate of 0.05.
Results: The results of our analysis are as follows:
@@
    No rescaling: Among all the demographic variables, the following variables have a significant association with the similarity to the target image when testing with ANOVA: computer programming frequency, self-reported programming ability, outlook towards generative AI, age, gender, generative AI use, education, imagery writing skill, and self-reported occupational skills of critical thinking, active listening and quality control.
@@
    When we fit a linear model to evaluate directionality, we find that those who use critical thinking as job skills, and report little usage of generative AI or imagery writing skill are, on average, better at the task. Conversely, we find that older people, men, those with a more positive outlook regarding generative AI, those with self-reported occupational skill of quality control, and those who programmed more frequently performed worse on our experiment task.
@@
    When conducting the non-parametric Kruskal-Wallis test, we found fewer significant variables. In this case, the statistically significant relationships with performance are: self-reported programming frequency, outlook towards generative AI, self-reported programming skill, gender and age.
    Z-score rescaling: Among all the demographic variables, the following variables have a significant association with the similarity to the target image when testing with ANOVA: computer programming frequency, self-reported programming ability and frequency, outlook towards generative AI, age, gender, generative AI use, education, imagery writing skill, and self-reported occupational skills of critical thinking, quality control, technology design, social perceptiveness and troubleshooting.
@@
    When we fit a linear model to evaluate directionality, we find that those who use critical thinking and troubleshooting as job skills, and report little usage of generative AI, little imagery writing skill, some programming skill, and some instruction writing skill are, on average, better at the task. Conversely, we find that older people, men, those with a more positive outlook regarding generative AI, those with a graduate degree, those with self-reported occupational skill of quality control and technology design, and those who programmed more frequently performed worse on our experiment task.
    When conducting the non-parametric Kruskal-Wallis test, we found fewer significant variables. In this case, the statistically significant relationships with performance are: self-reported programming frequency, outlook towards generative AI, self-reported programming skill, gender and age.
@@
    Percentile rank rescaling: Among all the demographic variables, the following variables have a significant association with the similarity to the target image when testing with ANOVA: computer programming frequency, self-reported programming ability and frequency, outlook towards generative AI, age, gender, generative AI use, and self-reported occupational skill of critical thinking.
    When we fit a linear model to evaluate directionality, we find that those who use critical thinking as a job skill, and report little usage of generative AI, and some programming skill are, on average, better at the task. Conversely, we find that older people, men, those with a more positive outlook regarding generative AI, those who programmed more frequently performed worse on our experiment task.
    When conducting the non-parametric Kruskal-Wallis test, we found fewer significant variables. In this case, the statistically significant relationships with performance are: self-reported programming frequency, outlook towards generative AI, self-reported programming skill, gender and age.
@@
    No rescaling: The following variables have a significant association with the similarity to the target image when testing with ANOVA: computer programming frequency, self-reported programming ability, outlook towards generative AI, age, gender, generative AI use, imagery and instructional writing skill, and self-reported occupational skills of critical thinking, learning strategies, technology design and quality control.
@@
    When we fit a linear model to evaluate directionality, we find that those who use critical thinking and social perceptiveness as job skills, and report little usage of generative AI, report little imagery writing or programming skills are, on average, better at the task. Conversely, we find that older people, men, those with a more positive outlook regarding generative AI, and those who programmed more frequently performed worse on our experiment task.
@@
    When conducting the non-parametric Kruskal-Wallis test, the following variables had a statistically significant relationships with performance: self-reported programming frequency, outlook towards generative AI, self-reported imagery writing skill, gender, age and learning strategies occupational skill.
    Z-score rescaling: The following variables have a significant association with the similarity to the target image when testing with ANOVA: computer programming frequency, self-reported programming ability, outlook towards generative AI, age, gender, generative AI use, imagery and instructional writing skills, education, and self-reported occupational skills of critical thinking, social perceptiveness, learning strategies, technology design and quality control.
    When we fit a linear model to evaluate directionality, we find that those who use critical thinking and social perceptiveness as job skills, and report little usage of generative AI, report little imagery writing or programming skills are, on average, better at the task. Conversely, we find that older people, men, those with a more positive outlook regarding generative AI, those who programmed more frequently, those with graduate degrees, and those with technology design, quality control and learning strategies as occupational skills performed worse on our experiment task.
@@
    When conducting the non-parametric Kruskal-Wallis test, the following variables had a statistically significant relationships with performance: self-reported programming frequency, outlook towards generative AI, self-reported imagery writing skill, gender, age and learning strategies, social perceptiveness and technology design occupational skills.
@@
    Percentile rank rescaling: The following variables have a significant association with the similarity to the target image when testing with ANOVA: computer programming frequency, self-reported programming ability, outlook towards generative AI, age, gender, generative AI use, imagery writing skill, education, and self-reported occupational skills of social perceptiveness, learning strategies, and technology design.
    When we fit a linear model to evaluate directionality, we find that those who use social perceptiveness as a job skill, and report little usage of generative AI, report little imagery writing or programming skills are, on average, better at the task. Conversely, we find that older people, men, those with a more positive outlook regarding generative AI, those who programmed more frequently, and those with technology design and learning strategies as occupational skills performed worse on our experiment task.
@@
    When conducting the non-parametric Kruskal-Wallis test, the following variables had a statistically significant relationships with performance: self-reported programming frequency, generative AI use, outlook towards generative AI, self-reported imagery writing skill, gender, age and learning strategies, social perceptiveness and technology design occupational skills.
@@
Analysis approach:
To investigate this hypothesis, we pre-registered evaluating how exploration/exploitation in the prompting space is related to performance. Hence, most of the pre-registered independent variables measured the level of similarity (i.e., exploitation) across a user’s prompts. We refer to participants being more “exploitative” if they wrote prompts that are more similar to their previous prompts and more “exploratory” if they wrote prompts that deviated more from their previous prompts. As described below, we operationalized “similarity” in a variety of ways.
@@
    Positively associated with exploitation: average token sort ratio compared to the previous prompt, average cosine similarity between the embeddings of a given prompt with the previous prompt, fraction of times a prompt contains the previous prompt as an exact substring
@@
    Negatively association with exploitation: the variance of the prompt embedding, the number of topical transitions
We also pre-registered measuring each participant’s average prompt length. These variables were measured per each user and across their first 10 attempts. Their association with performance, measured in terms of each of our eight outcome variables, was estimated in a linear model with DALL-E version fixed effects.
We also conducted a similar analysis at the iteration level to answer the following question: how is exploration/exploitation associated with subsequent performance in the next attempt, and is the strength of this relationship mediated by the quality of the previous attempt? To answer this question, first, we divided user-iteration observations into 6 equal-sized brackets by performance in the previous iteration. The bracketing allows us to explore heterogeneity in prompting behavior by the quality of previous attempts. We then estimated the effect of textual similarity to the previous prompt on the quality of the next attempt within each bracket. Our estimates adjusted for other covariates by matching user-iteration observations on the target image, DALL-E model, iteration and the exact quality of the previous attempt (Sävje et al., 2021).
Results: The results of our analysis are as follows:

### v5

@@ E.2 Task Performance and ATEs
The top pane of Figure 1B compares the average performance across models and attempt numbers (also referred to as iterations). It displays the average cosine similarity score stratified by the reference image. A notable feature in this figure is the performance dip during the second recreation attempt across both treatment arms. This is likely due to participants’ initial misunderstanding of the model’s “memoryless” nature. Participants failed to recognize that context from previous prompts was not carried over to new iterations. We observed numerous prompts in the second iteration across users that explicitly referenced the first prompt, a behavior that rarely occurred in subsequent attempts. However, from the third prompt onward, participants appeared to grasp the independence of each attempt, as evidenced by a marked decrease in cross-prompt references and a corresponding rebound in performance.
@@
The bottom pane of Figure 1B shows the average treatment effect (ATE) per iteration, which is the difference between the stratified averages of DALL-E 3 and DALL-E 2 in the top pane. [^10] In Section E, when we refer to “DALL-E 3”, we mean “DALL-E 3 (Verbatim)” unless otherwise specified. To test the widening impact of using DALL-E 3 on performance relative to DALL-E 2, we run the following fixed effects linear model with participant-level ($i$) clustered standard errors where iteration is treated as a numeric variable:
$Y_{i,n,t}=$ $\beta_{0}+\beta_{1}\,\text{iteration}+\beta_{2}\,\mathbb{I}[\text{dalleVersion = 3}]_{i}$ (19)
@@
$+\beta_{3}\,\text{iteration}\times\mathbb{I}[\text{dalleVersion =3}]_{i}+\gamma_{t}+\epsilon_{i,n,t}$
@@ E.3 Prompt Characteristics
Figure 1C compares the prompt length and prompt similarity of the two models. To generate these results, we first remove any prompt that does not constitute a good-faith attempt according to the sample construction procedure detailed in Section C.C.4. The prompt length is the average number of words per model and iteration stratified by the reference image. The prompt similarity is the average cosine similarity between all consecutive pairs of user prompts, which are both determined to be valid attempts, stratified by the reference image (see Section C.C.5.3 for details on similarity calculations).
The color scale in Figure 1C shows the stratified average similarity to the previous prompt across all users per each model. We find that superior model users write prompts that, on average, have $\beta=0.0184$ higher in cosine similarity to their previous prompts using cluster robust standard errors at the participant level ($p=0.0236$).
@@
Comparing the aggregate similarity of all attempts made by a given participant, we also find the prompts from DALL-E 3 participants were, on average, more similar than the prompts of the inferior model participants. For this analysis, we use the dispersion around the centroid in the prompt embedding space, explained in Section C.C.5.3, as the dependent variable. When we average across all participants by model, we find that the average distance of prompts written by superior model participants to their centroid is $\beta=0.0191$ smaller than inferior users ($p=0.0083$). Standard errors are cluster-robust at the participant level.
E.4 ATE Decomposition
@@
Figure 2 in the main text decomposes the ATE into the model and prompting effects. This decomposition is conceptually similar to a simple mediation analysis, with an important difference being that we can observe counterfactual outcomes (e.g., prompting the superior model as if it is the inferior model). This is not typically the case in mediation analysis, and makes causal identification rely on fewer assumptions.
To obtain counterfactual outcomes, we fed or “replayed” the participant prompts when interacting with one model (e.g., inferior) on another model (e.g., superior). The notation (prompt, model) specifies which treatment arm the prompts were written under and which model was used in the replay. For example, (2,3) indicates replaying prompts written under DALL-E 2 on DALL-E 3. To be clear, (2,2) and (3,3) correspond to the original observed treatment arms, while (2,3) and (3,2) are the counterfactual outcomes of interest. [^11] To avoid problems with model drift, we regenerated images for all four possible combinations at the same time and used these images for all analyses in the main text.
@@
The left-most point in Figure 2 corresponds to the average CLIP cosine similarity to the target image of (2,2). To make the interpretation of the results clearer, we have subtracted this quantity from all average quality scores and added a dashed line throughout. The second point from the left corresponds to average similarity to the target of (2,3), the third point from the left to (3,3), and the rightmost points to (3,2). All average similarity scores are stratified by iteration and reference image, and the standard errors are bootstrapped and cluster-robust at the participant level.
The model effect, as shown by the red braces in Figure 2, corresponds to the average increase in quality of (2,3) relative to (2,2). In the terminology of mediation analysis, the model effect would be referred to as the direct effect. The prompting effect, as shown by the blue braces in Figure 2, corresponds to the average increase in quality of (3,3) relative to (2,3). In the terminology of mediation analysis, the prompting effect would be referred to as the indirect effect. We can also test the difference in average quality between (3,3) and (3,2), as well as the difference between (3,2) and (2,2). Both of these differences are visible in Figure 2; the second is small and not statistically significant.
@@
The standard errors in Figure 2 correspond to the uncertainty around the estimated average score for each of the four replay conditions. These uncertainty estimates are insufficient for exact inference on the direct, indirect, and treatment effects. The statistics and significance values reported in the main text, which correspond to such effects (i.e., the difference between average estimates in two conditions) are obtained using a two-way (iteration and target image) fixed effect model with the effect type as the main independent variable:
@@
E.5 Heterogeneity by Skill
Table 1 in the main text demonstrates how total, model, and prompting effects vary across different user skill levels. The total effect compares the Cosine similarity of outputs from DALL-E 3 users replayed on DALL-E 3 to those from DALL-E 2 users replayed on DALL-E 2. The model effect compares outputs from DALL-E 2 users replayed on DALL-E 3 against the same users replayed on DALL-E 2. The prompting effect compares outputs from DALL-E 3 users replayed on DALL-E 3 to those from DALL-E 2 users replayed on DALL-E 3.
To assess user skill, users within each target image, iteration and replay scenario are divided into 50 equally sized brackets based on their performance, assigning each bracket a rank corresponding to performance deciles. For instance, when evaluating DALL-E 2 users replayed on DALL-E 2, each user’s skill level is determined by their percentile rank among all DALL-E 2 users replayed on DALL-E 2. Conversely, when the same user is replayed on DALL-E 3 (to estimate the model effect or to serve as a baseline in estimating the prompting effect), their performance decile is determined by their rank relative to all DALL-E 2 users replayed on DALL-E 3, per each iteration and target image.
After assigning performance deciles, we estimate the following linear model with two-way fixed effects (iteration and target image) for each effect type (model, prompting, and total):
$Y_{i,n,t}=$ $\beta_{0}+\beta_{1}\text{effect}+\beta_{2}\text{effect}\times\text{Performance Decile}_{i,n,t}\,+$ (21)
$\alpha_{n}+\gamma_{t}+\epsilon_{i,n,t}$
where $\beta_{1}$ represents the coefficient associated with the effect type being analyzed (model, prompting, or total), and $\beta_{2}$ captures the interaction between the effect type and user skill level. User skill is measured as the user’s rank (binned) within the same iteration, target image, and replay scenario. Standard errors are calculated using cluster-robust methods at the participant level.
This methodology closely follows the procedure described earlier in section E.4, with the additional step of incorporating an interaction term to evaluate skill-level heterogeneity.
Appendix F Robustness Checks
This section provides additional analyses that confirm the robustness of our main findings using alternative metrics and statistical approaches.
F.1 DreamSim-based Analysis
As discussed in Section C, we repeated all main-text analyses using DreamSim as an alternative image similarity metric. The results, presented below, demonstrate that our findings are not dependent on the specific similarity measure used.
F.1.1 Overall ATEs
@@
F.1.2 Figure 1
We reran the regression in Section E with $Y_{i,n,t}$ representing the DreamSim outcome instead of cosine similarity:
$Y_{i,n,t}=$ $\beta_{0}+\beta_{1}\,\text{iteration}+\beta_{2}\,\mathbb{I}[\text{dalleVersion = 3}]_{i}$ (22)
$+\beta_{3}\,\text{iteration}\times\mathbb{I}[\text{dalleVersion =3}]_{i}+\gamma_{t}+\epsilon_{i,n,t}$
@@ F.1.3 Figure 2
F.2 Z-score-based Analysis
As we discuss in Section D.2, we repeated all main-text analyses using the within-image-attempt Z-score of CLIP-cosine similarity to account for variation between images and attempts. When comparing across attempts, the Z-score was computed within each image, as mentioned in Section D.2. The results remain consistent throughout, and in some cases, the differences between the superior and inferior model are even more pronounced than when using raw cosine similarity.
F.2.1 Overall ATEs
@@ F.2.2 Figure 1
$\text{ZScore}_{i,n}=$ $\beta_{0}+\beta_{1}\,\text{iteration}+\beta_{2}\,\mathbb{I}[\text{dalleVersion = 3}]_{i}$ (23)
$+\beta_{3}\,\text{iteration}\times\mathbb{I}[\text{dalleVersion =3}]_{i}+\epsilon_{i,n}$
@@ F.2.3 Figure 2
Appendix G Pre-registration
Prior to conducting our experiment, we developed and pre-registered a comprehensive set of hypotheses and analyses. This pre-registration is deposited at OSF at the following URL: URL removed to maintain anonymity. While the main text focuses on a subset of our pre-registered analyses that constitute the most important and timely findings, this appendix provides a comprehensive overview of all analyses specified in our pre-registration.
In our pre-registration, we outlined plans to conduct each analysis using six [^12] See section G.2 on deviations from the pre-registration for the remaining two pre-registered outcome variables. possible outcome variables, all representing different transformations of the same underlying data. These include CLIP embedding cosine similarity and DreamSim, each with three rescaling methods:
  • No rescaling: The outcome variable used in its original form.
  • Z-score rescaling: The outcome variable transformed into a Z-score following the procedure detailed in Section D.2.
  • Percentile rank rescaling: The outcome variable converted to a percentile rank relative to all other prompts submitted for the same target image.
While we are still in the process of completing all analyses with each of these outcome variables, our preliminary results suggest that our findings are robust across these different transformations. As additional analyses are completed, we will update this appendix. Full results, data, and replication code available online upon publication.
G.1 Hypotheses and Results
Below, we present each hypothesis exactly as stated in our pre-registration document, followed by a summary of our findings. This comprehensive review demonstrates the robustness of the key results highlighted in the main text, while also providing additional insights into the relationship between model capabilities, user behavior, and task performance.
@@ G.1 Hypotheses and Results
Analysis approach: For this hypothesis, we conducted multiple ANOVA tests, one for each demographic variable, with the relevant performance measure as the dependent variable and the model (DALL-E version) as a covariate. As a robustness check, we repeated these analyses using the non-parametric Kruskal-Wallis U test. To account for multiple comparisons, we applied the Benjamini-Hochberg adjustment with a false discovery rate of 0.05.
Results: Our analyses revealed several consistent demographic patterns in prompt engineering ability across different outcome measures:
@@
    • No rescaling: Using ANOVA, we found significant associations between performance and: computer programming frequency, self-reported programming ability, outlook towards generative AI, age, gender, generative AI use, education, imagery writing skill, and self-reported occupational skills (particularly critical thinking, active listening, and quality control).
@@
      In linear models examining directionality, we found that participants who reported critical thinking as a job skill and little usage of generative AI or imagery writing skill performed better, on average. Conversely, older participants, men, those with more positive outlooks regarding generative AI, those reporting quality control as an occupational skill, and frequent programmers performed worse on our task.
@@
      The more conservative Kruskal-Wallis tests identified fewer significant variables: self-reported programming frequency, outlook towards generative AI, self-reported programming skill, gender, and age.
    • Z-score rescaling: ANOVA tests revealed significant associations with similar demographic factors as the unscaled measure, plus additional significant associations with self-reported occupational skills in technology design, social perceptiveness, and troubleshooting.
@@
      Linear models showed that participants reporting critical thinking and troubleshooting as job skills, little generative AI use, little imagery writing skill, some programming skill, and some instruction writing skill performed better. Worse performance was associated with older participants, men, positive generative AI outlook, graduate degrees, self-reported quality control and technology design skills, and frequent programming experience.
      Kruskal-Wallis tests again identified a smaller set of significant predictors: self-reported programming frequency, generative AI outlook, programming skill, gender, and age.
@@
    • Percentile rank rescaling: ANOVA identified significant associations between performance and computer programming frequency, self-reported programming ability and frequency, generative AI outlook, age, gender, generative AI use, and critical thinking as an occupational skill.
      Linear models showed better performance among those reporting critical thinking as a job skill, little generative AI use, and some programming skill. Worse performance was associated with older participants, men, positive generative AI outlook, and frequent programming experience.
      Kruskal-Wallis tests revealed the same set of significant predictors as with other scalings: programming frequency, generative AI outlook, programming skill, gender, and age.
@@
    • No rescaling: ANOVA tests identified significant associations between performance and: programming frequency, self-reported programming ability, generative AI outlook, age, gender, generative AI use, imagery and instructional writing skills, and several occupational skills (critical thinking, learning strategies, technology design, and quality control).
@@
      Linear models showed better performance among those reporting critical thinking and social perceptiveness as job skills, little generative AI use, and little imagery writing or programming skills. Worse performance was associated with older participants, men, positive generative AI outlook, and frequent programming experience.
@@
      Kruskal-Wallis tests identified significant relationships with: programming frequency, generative AI outlook, imagery writing skill, gender, age, and learning strategies as an occupational skill.
    • Z-score rescaling: Using ANOVA, significant associations included all factors identified in the unscaled analysis, plus education and additional occupational skills (social perceptiveness).
      Linear model directional findings were consistent with the unscaled measure, with additional negative associations for graduate degrees and self-reported technology design, quality control, and learning strategies skills.
@@
      Kruskal-Wallis tests showed significant relationships with: programming frequency, generative AI outlook, imagery writing skill, gender, age, and occupational skills in learning strategies, social perceptiveness, and technology design.
@@
    • Percentile rank rescaling: ANOVA identified significant associations with: programming frequency, programming ability, generative AI outlook, age, gender, generative AI use, imagery writing skill, education, and occupational skills in social perceptiveness, learning strategies, and technology design.
      Linear models showed better performance among participants reporting social perceptiveness as a job skill, little generative AI use, and little imagery writing or programming skills. Worse performance was associated with older participants, men, positive generative AI outlook, frequent programming experience, and self-reported technology design and learning strategies skills.
@@
      Kruskal-Wallis tests revealed significant relationships with: programming frequency, generative AI use, generative AI outlook, imagery writing skill, gender, age, and occupational skills in learning strategies, social perceptiveness, and technology design.
Across these analyses, several consistent patterns emerge. Interestingly, we found that frequent programming experience was associated with worse task performance, contrary to what might be expected. Men and older participants also performed worse on average. Those reporting critical thinking or social perceptiveness as job skills tended to perform better, while participants with positive outlooks toward generative AI and higher generative AI usage performed worse. These patterns were generally consistent across different outcome measures and scaling approaches, suggesting robust demographic differences in prompt engineering ability.
@@
Analysis approach: We investigated this hypothesis by examining the relationship between exploration/exploitation strategies in the prompting space and performance outcomes. We characterized participants as more ”exploitative” if they wrote prompts that were similar to their previous prompts, and more ”exploratory” if their prompts exhibited greater deviation from previous attempts. We operationalized these concepts using several metrics:
@@
    Measures positively associated with exploitation:
    • Average token sort ratio compared to the previous prompt
    • Average cosine similarity between the embeddings of consecutive prompts
    • Fraction of times a prompt contains the previous prompt as an exact substring
@@
    Measures negatively associated with exploitation:
    • Variance of the prompt embedding
    • Number of topical transitions
We also measured each participant’s average prompt length. These variables were calculated for each user across their first 10 attempts, and their association with performance was estimated using linear models with DALL-E version fixed effects.
To examine how exploration/exploitation strategies influence performance across successive attempts, we conducted an additional analysis at the iteration level. We divided user-iteration observations into 6 equal-sized brackets based on performance in the previous iteration, allowing us to explore how prompting behavior might vary depending on the quality of previous attempts. We then estimated the effect of textual similarity to the previous prompt on the quality of the next attempt within each bracket. Our estimates adjusted for other covariates by matching user-iteration observations on target image, DALL-E model, iteration number, and exact quality of the previous attempt (Sävje et al. 2021).
Results: Our analyses revealed consistent patterns across different outcome measures: